100 ml of x molar H2O2 solution gives 3 lit. O2 gas under conditions when 1 mol occupies 24.lit. value of x

Grade:12th pass

1 Answers

Khimraj
5 years ago
 
from equation, 24 L of o2 comes from 2 moles of h2o2.
                      3 L of o2 comes from ¼ moles of h2o2
  100 mL=10-1 L. therefore x/10=1/4. assuming molarity is x.
          x=2.5M. please approve if its right